# Q-PETR: Quant-aware Position Embedding Transformation for Multi-View 3D Object Detection

### v1

@@ 4 Experiment
4.1 Experimental Setup
@@ 4.2 Ablation Study
Proof of Position Encoding Equivalence We conducted experiments to verify whether the proposed QD-aware lidar-ray PE enhances floating-point performance over the original camera-ray PE. As shown in Tab. 2, QD-aware lidar-ray PE provides performance improvements. On PETR, it slightly increases mAP by 0.07 but significantly boosts NDS and mATE by 1.09 and 1.67, respectively. For Stream-PETR, our method yields substantial and balanced enhancements, with increases of 0.94 in mAP, 0.46 in NDS, and 0.22 in mATE.
@@
| PETR | Camera-ray PE | 31.42(0.00) | 36.11(0.00) | 84.19(0.00) |
| QD-aware lidar-ray PE | 31.49(0.07) | 37.20(1.09) | 82.52(1.67) | |
| Stream -PETR | Camera-ray PE | 48.19(0.00) | 57.11(0.00) | 60.99(0.00) |
| QD-aware lidar-ray PE | 49.13(0.94) | 57.57(0.46) | 60.77(0.22) | |
Effect of Anchor Embedding Quantity. The QD-aware lidar-ray PE uses three anchor embeddings per axis, obtained through linear interpolation. Experiments (Tab. 3) demonstrate that setting the number of anchor embeddings to 3 achieves the highest NDS and mAP scores. Adjusting this number either up or down results in lower performance, confirming that 3 is the optimal choice.
@@
Quantization Performance of Different Position Encodings To experimentally demonstrate the superior quantization performance of our proposed QD-aware lidar-ray PE, we focus solely on quantizing the positional encoding, keeping all other modules in floating-point computation. Detailed results are shown in Tab.4. The original camera-ray configuration loses up to 11.97% in mAP and 5.04% in NDS, whereas our QD-aware lidar-ray PE experiences minimal losses of only 1.42% in mAP and 1.15% in NDS. Figure8 further supports this finding; compared to the distribution in Fig. 3, the distribution of our QD-aware lidar-ray PE aligns more closely with that of image features, retaining sufficient useful information.
@@
| QD-aware lidar-ray PE | 56.81 | 47.65 | 61.53 | |
Impact of Different Scaled Dot Product Quantization Strategy. To validate our novel scaled dot-product quantization strategy—which searches for the optimal scaling factor by minimizing softmax output error—we focus solely on quantizing the softmax inputs while keeping all other modules in floating-point computation. As shown in Tab. 5, the original quantization strategy results in significant losses of 40% in NDS and 50% in mAP. In contrast, our ”quant after stabilization” approach greatly improves performance. An ablation study on the maximum candidate truncation range reveals that setting yields optimal quantization performance with nearly no loss. Performance deteriorates when due to increased truncation of feature information, while values of exceeding 20 offer no additional benefits. Therefore, setting is sufficient to achieve the best performance.
@@
| quant before stabilization | - | 25.31 | 13.79 | 107.94 |
| quant after stabilization | N = 10 | 3.45 | 1.23 | 150.34 |
@@
| Backbone | Input Size | Feature | Mode | mAP | NDS | mATE | mASE | mAOE | mAVE | mAAE | |
| R50 | 5121408 | c5 | PETR | FP32 | 31.42(0.00) | 36.11(0.00) | 84.19(0.00) | 28.42(0.00) | 60.69(0.00) | 99.08(0.00) | 23.58(0.00) |
| PTQ | 13.79[56.11] | 25.31[29.91] | 107.94[28.21] | 31.47[10.73] | 75.22[23.94] | 83.71[15.51] | 25.45[7.93] | ||||
| Q-PETR | FP32 | 31.49(0.07) | 37.20(1.09) | 82.52(1.67) | 27.88(0.54) | 59.91(0.78) | 91.74(7.34) | 23.45(0.13) | |||
| PTQ | 31.34[0.47] | 37.17[0.82] | 82.61[0.65] | 27.93[0.17] | 60.00[0.15] | 91.79[0.00] | 23.45[0.00] | ||||
| R50 | 5121408 | p4 | PETR | FP32 | 32.60(0.00) | 37.16(0.00) | 82.63(0.00) | 27.96(0.00) | 61.06(0.00) | 95.81(0.00) | 23.91(0.00) |
| PTQ | 12.97[60.21] | 24.75[33.39] | 108.28[31.04] | 31.76[13.59] | 79.57[30.31] | 78.90[17.65] | 27.14[13.51] | ||||
| Q-PETR | FP32 | 32.69(0.09) | 38.03(0.87) | 80.58(2.05) | 27.89(0.07) | 59.43(0.63) | 92.69(3.12) | 22.55(1.36) | |||
| PTQ | 32.40[0.88] | 37.72[0.82] | 81.11[0.65] | 27.92[0.10] | 60.02[0.97] | 92.76[0.00] | 22.59[0.18] | ||||
| R101 | 5121408 | p4 | PETR | FP32 | 34.40(0.00) | 38.62(0.00) | 80.67(0.00) | 28.03(0.00) | 57.13(0.00) | 95.74(0.00) | 24.20(0.00) |
| PTQ | 13.53[60.67] | 23.84[38.27] | 111.04[37.65] | 31.27[11.56] | 78.94[38.18] | 92.92[2.95] | 26.14[8.02] | ||||
| Q-PETR | FP32 | 34.72(0.32) | 39.68(1.06) | 79.40(1.27) | 27.92(0.11) | 53.90(3.23) | 92.99(2.75) | 22.59(1.61) | |||
| PTQ | 34.41[0.89] | 39.08[1.51] | 80.98[1.98] | 28.00[0.28] | 54.41[0.94] | 92.62[0.39] | 22.70[0.47] | ||||
| V2-99 | 320800 | p4 | PETR | FP32 | 38.01(0.00) | 42.56(0.00) | 75.79(0.00) | 26.84(0.00) | 50.58(0.00) | 90.13(0.00) | 21.07(0.00) |
| PTQ | 10.46[72.48] | 23.64[44.45] | 112.41[36.21] | 33.00[22.95] | 85.96[69.95] | 71.83[20.30] | 25.12[19.22] | ||||
| Q-PETR | FP32 | 38.43(0.42) | 43.80(1.24) | 74.79(1.00) | 27.29(0.45) | 49.76(0.82) | 82.11(8.02) | 20.15(0.92) | |||
| PTQ | 37.93[1.30] | 43.17[1.44] | 75.50[0.94] | 27.78[1.79] | 50.07[0.62] | 82.41[0.36] | 20.38[1.14] | ||||
| V2-99 | 6401600 | p4 | PETR | FP32 | 40.66(0.00) | 46.05(0.00) | 71.76(0.00) | 27.07(0.00) | 42.23(0.00) | 80.68(0.00) | 21.06(0.00) |
| PTQ | 6.40[84.63] | 20.98[54.55] | 117.38[33.22] | 34.85[25.92] | 83.38[97.44] | 76.83[4.79] | 27.10[28.67] | ||||
| Q-PETR | FP32 | 41.35(0.69) | 45.99(0.06) | 72.18(0.42) | 26.91(0.15) | 45.05(2.82) | 82.03(2.35) | 20.67(0.39) | |||
| PTQ | 40.95[0.96] | 45.64[1.09] | 73.40[1.69] | 27.05[0.52] | 45.61[1.24] | 82.17[0.17] | 20.68[0.00] | ||||
| V2-99 | 320800 | p4 | StreamPETR | FP32 | 48.19(0.00) | 57.11(0.00) | 60.99(0.00) | 25.58(0.00) | 37.54(0.00) | 26.28(0.00) | 19.43(0.00) |
| PTQ | 18.52(61.19) | 36.47(36.11) | 76.39(25.25) | 31.44 (22.90) | 47.03 (25.27) | 30.22 14.99 | 20.99 (8.03) | ||||
| Q-StreamPETR | FP32 | 49.13(0.94) | 57.57(0.46) | 60.77(0.22) | 26.14(0.56) | 39.05(1.49) | 24.81(1.47) | 19.15(0.28) | |||
| PTQ | 48.21 (1.87) | 56.33 (2.15) | 63.00 (3.66) | 26.35 (0.80) | 39.17 (0.31) | 24.90 (0.36) | 19.19 (0.21) | ||||
| V2-99 | 6401600 | p4 | StreamPETR | FP32 | 49.51(0.00) | 58.03(0.00) | 60.10(0.00) | 26.07(0.00) | 35.65(0.00) | 25.91(0.00) | 19.60(0.00) |
| PTQ | 18.72(62.19) | 35.66(38.54) | 74.32(23.66) | 30.39 (16.76) | 41.49 (16.38) | 30.53 17.83 | 20.82 (6.22) | ||||
| Q-StreamPETR | FP32 | 50.48(0.93) | 58.61(0.58) | 58.78(0.32) | 26.16(0.09) | 37.05(1.40) | 25.69(0.22) | 18.59(1.01) | |||
| PTQ | 49.44 (2.06) | 57.94 (1.24) | 60.30 (2.52) | 26.55 (1.49) | 37.07 (0.00) | 25.87 (0.31) | 18.59 (0.00) | ||||
@@
4.3 Validation on PETR-series Methods
We evaluate the effectiveness of our proposed method on various PETR-series models from both floating-point and quantized performance perspectives, specifically considering single-frame PETR and temporal multi-frame StreamPETR models.
First, we analyze changes in floating-point performance (values in parentheses in Tab. 6). In single-frame PETR models, mAP and NDS generally improve across configurations, except for a slight decrease of 0.06 in NDS when using V2-99’s P4 feature with 640×1600 resolution images. mAP increases range from 0.07 to 0.69, while NDS shows significant gains in all cases, ranging from 0.87 to 1.24. For temporal multi-frame StreamPETR models, both mAP and NDS consistently improve, with mAP gains of 0.93 and 0.94, and modest NDS increases of 0.46 and 0.58. Notably, NDS improvements in temporal methods are smaller than in single-frame methods, mainly due to performance degradation in mASE and mAOE, suggesting that our QD-aware lidar-ray PE may not optimally capture scale and orientation information in temporal models. We plan to investigate this further in future work. Overall, QPETR shows significant improvements in most configuration metrics, demonstrating that our method surpasses the original PETR models in floating-point performance.
Next, we analyze the quantized performance improvements (values in square brackets in Table 6). In single-frame PETR models, mAP and NDS drops are kept below 1% using our QD-aware lidar-ray PE and smoothing techniques. In temporal multi-frame StreamPETR models, mAP and NDS drops remain within 2.5%, likely due to accumulated quantization errors during temporal fusion. Overall, quantized QPETR models maintain high performance with minimal drops in both settings, demonstrating the effectiveness of our quantization strategies in preserving accuracy while reducing computational and memory requirements. We intend to further mitigate quantization errors in temporal models through enhanced error correction or advanced quantization methods.
5 Limitations
Although our method incurs almost no quantization accuracy loss, users need to replace the camera-ray in the original PETR series with our proposed QD-aware lidar-ray. The only drawback is that this requires retraining. However, from the perspective of quantization deployment, this retraining is beneficial, and the floating-point precision can even be improved.
@@ 6 Conclusion
In this paper, we address the significant performance drops of PETR models during quantization by identifying two main issues: the imbalance between positional encoding and image feature magnitudes, and uneven scalar dot-products in cross-attention. To resolve these, we introduce Q-PETR, a quantization-friendly positional encoding transformation that redesigns positional encoding and improves scalar dot-product quantization without sacrificing the original floating-point performance. Our experiments show that Q-PETR limits mAP and NDS drops to below 1% under standard 8-bit post-training quantization and even surpasses the original PETR in floating-point precision. Extensive tests across various PETR models demonstrate the method’s strong generalization and deployment suitability. Future work will explore quantization error propagation in multi-frame temporal models and develop advanced quantization strategies to further reduce precision loss in temporal settings.

### v2

@@ 4 Experiment
Detailed descriptions of benchmark, along with metric and further experimental details, are elaborated in Sup. B.
4.1 Validation on PETR-series Methods
We evaluate the effectiveness of our method on various PETR-series models from both FP and quantized performance perspectives, specifically considering single-frame PETR and temporal multi-frame StreamPETR models. Firstly, we analyze changes in floating-point performance (values in parentheses in Tab. 2). In single-frame PETR models, mAP and NDS generally improve across configurations, except for a slight decrease of 0.06 in NDS when using V2-99’s P4 feature with 640×1600 resolution images. mAP increases range from 0.07 to 0.69, while NDS shows significant gains in all cases, ranging from 0.87 to 1.24. For temporal multi-frame StreamPETR models, both mAP and NDS consistently improve, with mAP gains of 0.93 and 0.94, and modest NDS increases of 0.46 and 0.58. Notably, NDS improvements in temporal methods are smaller than in single-frame methods, mainly due to performance degradation in mASE and mAOE, suggesting that our QDPE may not optimally capture scale and orientation information in temporal models. We plan to investigate this further in future work. Overall, QPETR shows significant improvements in most configuration metrics, demonstrating that our method surpasses the original PETR models in floating-point performance. Secondly, we analyze the quantized performance improvements. In single-frame PETR models, mAP and NDS drops are kept below 1% using our QDPE and smoothing techniques. In temporal multi-frame StreamPETR models, mAP and NDS drops remain within 2.5%, likely due to accumulated quantization errors during temporal fusion. Overall, quantized QPETR models maintain high performance with minimal drops in both settings, demonstrating the effectiveness of our quantization strategies in preserving accuracy while reducing computational and memory requirements. We intend to further mitigate quantization errors in temporal models through enhanced error correction or advanced quantization methods.
@@ 4.2 Ablation Study
Proof of Position Encoding Equivalence. We conducted experiments to verify whether the proposed QDPE enhances floating-point performance over the original camera-ray PE. As shown in Tab. 3, QDPE provides performance improvements. On PETR, it slightly increases mAP by 0.07 but significantly boosts NDS and mATE by 1.09 and 1.67, respectively. For Stream-PETR, our method yields substantial and balanced enhancements, with increases of 0.94 in mAP, 0.46 in NDS, and 0.22 in mATE.
@@
| PETR | Camera-ray PE | 31.42 | 36.11 | 84.19 |
| QDPE | 31.49 | 37.20 | 82.52 | |
| Stream -PETR | Camera-ray PE | 48.19 | 57.11 | 60.99 |
| QDPE | 49.13 | 57.57 | 60.77 | |
Quantization Performance Evaluation. We evaluate the Camera-ray PE module on the nuScenes dataset under three configurations: FP32 Baseline (full precision as an upper bound), Standard 8-bit PTQ (per-tensor 8-bit post-training quantization), and PTQ4ViT (using the PTQ4ViT [55] method to boost accuracy). As shown in Table 4, retaining the Softmax input in full precision (“No”) yields higher mAP and NDS than when it is quantized (“Yes”), underscoring the importance of careful Softmax treatment.
| Method | Quant. Softmax Input | (PTQ) INT8 | (PTQ4ViT) INT8 | ||
| mAP | NDS | mAP | NDS | ||
| Camera- ray PE | FP32 | 31.42 | 36.11 | 31.42 | 36.11 |
| No | 24.90 | 32.10 | 27.10 | 33.60 | |
| Yes | 18.80 | 27.50 | 19.20 | 28.40 | |
Compare with QAT. Although QDPE requires retraining, its cost is comparable to that of QAT. We implement a distillation-based QAT (inspired by QD-BEV [57]) on the original Camera-ray PE. As shown in Table 5, even after 24 or 36 epochs, QAT yields lower mAP and NDS than QDPE with standard PTQ. This confirms that our amplitude-aware design not only maintains high floating-point performance but also achieves superior quantized accuracy.
| Model | 12 epochs | 24 epochs | 36 epochs | ||||
| mAP | NDS | mAP | NDS | mAP | NDS | ||
| Camera-ray PE | QAT (Distill) | 28.9 | 35.2 | 30.3 | 35.8 | 30.3 | 35.8 |
| QDPE | PTQ | 31.34 | 37.17 | 31.34 | 37.17 | 31.34 | 37.17 |
Effect of Anchor Embedding Quantity. The QDPE uses three anchor embeddings per axis, obtained through linear interpolation. Experiments (Tab. 6) demonstrate that setting the number of anchor embeddings to 3 achieves the highest NDS and mAP scores. Adjusting this number either up or down results in lower performance, confirming that 3 is the optimal choice.
@@
Quantization Performance of Different Position Encodings. To experimentally demonstrate the superior quantization performance of our proposed QDPE, we focus solely on quantizing the positional encoding, keeping all other modules in floating-point computation. Detailed results are shown in Tab. 7. The original camera-ray configuration loses up to 11.97% in mAP and 5.04% in NDS, whereas our QDPE experiences minimal losses of only 1.42% in mAP and 1.15% in NDS. Fig. 9 further supports this finding; compared to the distribution in Fig. 4, the distribution of our QDPE aligns more closely with that of image features, retaining sufficient useful information.
@@
| QDPE | 56.81 | 47.65 | 61.53 | |
Impact of Different Scaled Dot Product Quantization Strategy. To validate our novel scaled dot-product quantization strategy—which searches for the optimal scaling factor by minimizing softmax output error—we focus solely on quantizing the softmax inputs while keeping all other modules in floating-point computation. As shown in Tab. 8, the original quantization strategy results in significant losses of 40% in NDS and 50% in mAP. In contrast, our ”quant after stabilization” approach greatly improves performance. An ablation study on the maximum candidate truncation range reveals that setting yields optimal quantization performance with nearly no loss. Performance deteriorates when due to increased truncation of feature information, while values of exceeding 20 offer no additional benefits. Therefore, setting is sufficient to achieve the best performance. Additionally, in large language models, the attention inputs can reach extremely large values (see Fig. 10), and we have validated the effectiveness of our method in such scenarios as well (see Tab. 9).
@@
| quant before ns | - | 25.31 | 13.79 | 107.94 |
| quant after ns | N = 10 | 3.45 | 1.23 | 150.34 |
@@
| Model Name | qwen2.5-7b-instruct | deepseek-r1-distill-qwen-7b | ||
| wikitext2 | gsm8k | wikitext2 | gsm8k | |
| bfp16 | 7.46 | 80.21 | 25.04 | 85.97 |
| quant before ns | 10000+ | 0.3 | 10000+ | 0.1 |
| quant after ns(20) | 7.48 | 80.24 | 25.09 | 86.03 |
@@
Practical Hardware Resource Savings. Tab. 11 shows Q-PETR runs at 13.3 FPS (87% faster) and 1.9 GB memory (60% less) vs. PETR’s 7.1 FPS/4.8 GB, demonstrating significant speedup and resource efficiency.
| Method | Mode | FPS | CUDA memory (G) |
| PETR | fp32 | 7.1 | 4.8 |
| Q-PETR | INT8 | 13.3 | 1.9 |
@@ 5 Conclusion
In this paper, we address the significant performance drops of PETR models during quantization by identifying two main issues: the imbalance between positional encoding and image feature magnitudes, and uneven scalar dot-products in cross-attention. To resolve these, we introduce Q-PETR, a quantization-friendly positional encoding transformation that redesigns positional encoding and improves scalar dot-product quantization without sacrificing the original floating-point performance. We also propose DuLUT, a dual-table lookup mechanism for efficiently quantizing nonlinear functions, further enhancing deployment suitability on edge AI chips. Our experiments show that Q-PETR limits mAP and NDS drops to below 1% under standard 8-bit post-training quantization and even surpasses the original PETR in floating-point precision. Extensive tests across various PETR models demonstrate the method’s strong generalization and deployment suitability.
Supplementary
Appendix A Preliminaries
PETR
Appendix B Experimental Setup
Appendix C Theoretical Analysis of Magnitude Bounds in Position Encodings
C.1 Normalization Framework and Input Conditioning
To establish a unified analytical framework, we first formalize the spatial normalization process for various ray-based position encodings. Let denote the 3D coordinates within the perception range meters and meters. The normalized coordinates are computed as:
Noting that is clamped to within the range , the distribution ranges of the normalized sampled points in positional encodings are characterized as follows:
For the sampled point of Camera-Ray PE, denoted as , the distribution spans the unit cube, i.e., .
For the sampled points of LiDAR-Ray PE and QDPE, denoted as and respectively, the distributions are constrained to .
Here, the value is derived from the ratio , where corresponds to the fixed depth setting in the encoding process. This distinction highlights the inherent differences in spatial coverage and normalization strategies employed by these positional encodings.
C.2 Magnitude Propagation Analysis
C.2.1 Camera-Ray Position Encoding
As illustrated in Fig. 6 (a), the encoding pipeline consists of two critical stages:
Stage 1: Inverse Sigmoid Transformation
Empirical analysis reveals a maximum magnitude .
Stage 2: MLP Projection (Through Two Fully-Connected Layers)
where denotes the ReLU activation function. Let be the maximum weight magnitude. We derive the upper bound:
where and denote the input tensor channels for and , respectively.
C.2.2 LiDAR-Ray Position Encoding
Unlike Camera-Ray PE, the encoding process of LiDAR-Ray PE introduces sinusoidal modulation between the inverse sigmoid transformation and MLP projection, as shown in Fig. 6 (b). The magnitude propagation for LiDAR-Ray PE is as follows:
Stage 1: Inverse Sigmoid Transformation
Empirical analysis reveals a maximum magnitude .
Stage 2: Spectral Embedding
where denotes concatenation. This ensures:
Stage 3: MLP Projection (Following setting in Camera-Ray PE)
C.2.3 Ours QD-PE
The proposed encoding introduces anchor-based constraints, as depicted in Fig. 6 (c):
Stage 1: Anchor Interpolation (For Each Axis )
where denotes learnable anchor embeddings. Via Theorem C.1, the magnitude is constrain to:
Stage 2: MLP Projection
C.3 Comparative Magnitude Analysis
The derived bounds reveal fundamental differences in magnitude scaling:
This analysis demonstrates that QD-PE requires less quantization range than Camera-Ray PE.
C.4 Theoretical Guarantee of Magnitude Constraints
Theorem C.1 (Anchor Embedding Magnitude Bound).
Let be adjacent anchor embeddings with . For any point , its interpolated embedding satisfies:
Proof.
Let . The interpolated embedding becomes:
For any component :
Thus, holds for all dimensions. ∎
Through the application of regularization (e.g., L2 constraint) on the anchor embeddings during training, the magnitude of can be explicitly controlled. Empirically, we find that this value converges to approximately 0.8 in our experiments.
Appendix D More Ablation Study
D.1 Local Similarity of Position Encoding Features
Fig. 11 shows that QD-PE significantly outperforms 3D point PE and cameraray PE in local similarity of position encoding. Its similarity distribution appears more compact and concentrated, validating the method’s superiority in local spatial information modeling and its capability to precisely capture neighborhood spatial relationships around target pixels.
Appendix E Limitations
Although our method incurs almost no quantization accuracy loss, users need to replace the camera-ray in the original PETR series with our proposed QDPE. The only drawback is that this requires retraining. However, from the perspective of quantization deployment, this retraining is beneficial, and the floating-point precision can even be improved.